If output is 80,000, consumption is 40,000, investment is 20,000 and government spending is 10,000, is this country running a trade deficit or a trade surplus? Explain.
Answer #1

We use a simple formula Y = C + I + G + (EX - IM). In case of a deficit, the imports are more than the exports.

As per the given scenario above total output or Y is 80,000. Consumption is 40,000, investment is 20,000 and government spending is 10,000.

80,000 = 40,000 + 20,000 +10,000 + (Net exports).

80,000 = 70,000 + NE

NE = 80,000 - 70,000

NE = 10,000.

In this case, we have an export surplus. The country is running an export surplus of $10,000. (If the number was negative we would have a deficit. That means out consumption is more and output is less the gap is filled by imports. )